OpenStudy (anonymous):

Find an equation for the tangent to the curve at the given point. 11) f(x) = 4 x - x + 3, (16, 3)

OpenStudy (e.mccormick):

Use a derivative to find the slope pf the tangent line at that point. Then use point/slope formula to find the equation of the tangent line.

OpenStudy (e.mccormick):

of

OpenStudy (e.mccormick):

The given equation.

OpenStudy (anonymous):

f'(x)=3

OpenStudy (e.mccormick):

Yah... hehe. They gave you an easy one for this.

OpenStudy (anonymous):

now* what do i do

OpenStudy (e.mccormick):

Point/Slope formula. Remember that from Algebra?

OpenStudy (anonymous):

it may be x-x1=slope -x1-x??

OpenStudy (e.mccormick):

y and x.

OpenStudy (e.mccormick):

\(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

so how does the derivative fit into that?

OpenStudy (e.mccormick):

it is m.

OpenStudy (anonymous):

so m is 3? or do i need to figure out something

OpenStudy (e.mccormick):

Your \((x_1,y_1)\) is the point given, (16, 3). When you see the answer, look closely at the original question because it is linear.

OpenStudy (anonymous):

so it is y-3=3-16-x

OpenStudy (anonymous):

y=-10x

OpenStudy (e.mccormick):

y-3=3(x-16)

OpenStudy (anonymous):

so i solve that?

OpenStudy (anonymous):

y=6x-45

OpenStudy (e.mccormick):

Yes, but I am kind of wondering about the question you typed in. The point you gave does not fall on that line.

OpenStudy (e.mccormick):

6x?

OpenStudy (e.mccormick):

-45?

OpenStudy (anonymous):

ohh ok its missing something

OpenStudy (e.mccormick):

Actually, yah... 45 is right.

OpenStudy (anonymous):

it is f(x)=4 square root x -x+3

OpenStudy (e.mccormick):

\[y-3=3(x-16)\implies y-3=3x-48\implies y=3x-45\] OH! \(4 \sqrt{ x} -x+3\)?? That would be a completley different line!

OpenStudy (anonymous):

f(x)=4√x-x+3

OpenStudy (anonymous):

what to do?

OpenStudy (e.mccormick):

OK... We start over! Wrong derivative because we were working the wrong line.

OpenStudy (e.mccormick):

Slope should be negative with that line.... the root one.

OpenStudy (anonymous):

so i find the derivative?

OpenStudy (e.mccormick):

Yep!

OpenStudy (anonymous):

f'(x)+1=2/squareroot(x)

OpenStudy (e.mccormick):

You would not move over the 1.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

f'(x)+1=2/√x

OpenStudy (e.mccormick):

You wrote: \(f'(x)+1=\frac{2}{\sqrt{x}} \) But it is: \(f'(x)=\frac{2}{\sqrt{x}} - 1\)

OpenStudy (anonymous):

oh ok

OpenStudy (e.mccormick):

Now, this is more like how they normally do these. You need the slope at the point (16,3). So put the 16 into \(f'(x)\) to get the slope at that particular point.

OpenStudy (anonymous):

-0.5

OpenStudy (e.mccormick):

\(f'(x)=\frac{2}{\sqrt{x}} - 1\) \(f'(16)=\frac{2}{\sqrt{16}} - 1\) \(f'(16)=\frac{2}{4} - 1\) \(f'(16)=\frac{1}{2} - 1\) \(f'(16)=-\frac{1}{2}\) So slope of the tangent line at \((16,3)\) is \(-\frac{1}{2}\) Yep!

OpenStudy (anonymous):

y-3=-0.5(16-x)

OpenStudy (e.mccormick):

Now we are back to \(y-y_1=m(x-x_1)\) Not 16-x, but x-16.

OpenStudy (anonymous):

y=1/2x+11

OpenStudy (anonymous):

y=2.5x+11

OpenStudy (e.mccormick):

\(y-y_1=m(x-x_1)\) \(y-3=-\frac{1}{2}(x-16)\) \(y-3=-\frac{1}{2}x+8\) \(y=-\frac{1}{2}x+11\) And that is an equation for the tangent line at that point. Don;t foget to keep the - on the 1/2 x.

OpenStudy (anonymous):

ok thanks. but why isnt that half 2.5

OpenStudy (e.mccormick):

Why would 1/2 suddenly change to something else? The 3 from the other side is added, so it does not distribute to both terms.

OpenStudy (e.mccormick):

\(y-3=-\frac{1}{2}x+8\) \(y-3+3=-\frac{1}{2}x+8+3\) \(y=-\frac{1}{2}x+11\)

OpenStudy (anonymous):

oh i thought it did

OpenStudy (anonymous):

ok thanks

OpenStudy (e.mccormick):

Notice that even though this was a calculus problem, only one step was calculus. The majority was a bunch of algebra. That is how calculus is. Lots of algebra and trig. So if you want to be good at calculus, brush up on the prior stuff. It will help a ton!

OpenStudy (anonymous):

ok

OpenStudy (e.mccormick):

Have fun!